OpenStudy (meowlover17):

Given the following functions f(x) and g(x), solve (f + g)(5) and select the correct answer below: f(x) = 8x − 13 g(x) = x − 9

OpenStudy (meowlover17):

A.23 B.27 C.31 D.40

OpenStudy (meowlover17):

Please help

OpenStudy (anonymous):

add 8x - 13 + x - 9 then solve for x by substituting 5

OpenStudy (anonymous):

8(5) - 13 + (5) - 9

OpenStudy (meowlover17):

Thank you

OpenStudy (anonymous):

np x

OpenStudy (meowlover17):

So 23

OpenStudy (anonymous):

yes!!

OpenStudy (meowlover17):

Could you help me with one more?

OpenStudy (anonymous):

sure

OpenStudy (meowlover17):

Tablets are on sale for 15% off the original price (t), which can be expressed with the function p(t) = 0.85t. Local taxes are an additional 8% of the discounted price (p), which can be expressed with the function c(p) = 1.08p. Using this information, which of the following represents the final price of a tablet with the discount and taxes applied based on its original price?

OpenStudy (meowlover17):

c[p(t)] = 0.918t c(p) + p(t) = 1.93t c(p) ⋅ p(t) = 0.918pt t[c(p)] = 1.93p

OpenStudy (meowlover17):

I dont understand this at all.

OpenStudy (anonymous):

yeah i'm not sure

OpenStudy (meowlover17):

Another one?

OpenStudy (anonymous):

nengeni you know?

OpenStudy (meowlover17):

Given the following functions f(x) and g(x), solve f[g(6)] and select the correct answer below: f(x) = 6x + 12 g(x) = x − 8

OpenStudy (meowlover17):

@090909090909

OpenStudy (meowlover17):

−96 0 24 48

OpenStudy (anonymous):

okay this is super easy you got this since (g(6)) is in parentheses, you solve that first plug in 6 for x in the function g(x) g(x) = 6 - 8

OpenStudy (meowlover17):

Okay

OpenStudy (anonymous):

the answer you get from that is what you plug into the function f(x) as x

OpenStudy (meowlover17):

So i plug in -2 for X?

OpenStudy (meowlover17):

So my answer is 0 :D

OpenStudy (anonymous):

yes, f(x) = 6(-2) + 12

OpenStudy (anonymous):

yep! B

OpenStudy (meowlover17):

Thank you so much can you help me with a few more like that?

OpenStudy (anonymous):

sure post them

OpenStudy (meowlover17):

Alright awesome Given the following functions f(x) and g(x), solve (f ⋅ g)(3) and select the correct answer below: f(x) = 4x2 + 12 g(x) = x − 1

OpenStudy (meowlover17):

47 −47 96 −96

OpenStudy (anonymous):

is it \[4x^2 + 12\]

OpenStudy (meowlover17):

Yes

OpenStudy (anonymous):

aight. expanded out, this would be written as (4(3)^2 + 12)*(3 - 1)

OpenStudy (meowlover17):

So do i solve for that expanded form?2

OpenStudy (anonymous):

start with 4(3)^2 + 12

OpenStudy (anonymous):

follow order of operations, once you get an answer, multiply it by (3-1)

OpenStudy (meowlover17):

156?

OpenStudy (meowlover17):

312 0-0

OpenStudy (anonymous):

NO, NO

OpenStudy (meowlover17):

Oh wait sorry XD

OpenStudy (meowlover17):

i did 4*3^2

OpenStudy (anonymous):

order of operations : parentheses, exponents, multiplication

OpenStudy (anonymous):

yes so square 3 and then multiply

OpenStudy (meowlover17):

96

OpenStudy (anonymous):

yes

OpenStudy (meowlover17):

Thats my answer o-o

OpenStudy (anonymous):

yes

OpenStudy (meowlover17):

Awesome another one if thats fine by you

OpenStudy (anonymous):

yep

OpenStudy (meowlover17):

Given the functions k(x) = 5x − 8 and p(x) = x − 4, solve k[p(x)] and select the correct answer below.

OpenStudy (meowlover17):

k[p(x)] = 5x − 12 k[p(x)] = 5x − 28 k[p(x)] = 5x2 − 12 k[p(x)] = 5x2 − 28

OpenStudy (meowlover17):

its 5x^2

OpenStudy (anonymous):

what k[p(x)] means is: substitute the function p(x) as x in the function k(x)

OpenStudy (meowlover17):

0k

OpenStudy (meowlover17):

I dont understand o-o

OpenStudy (anonymous):

5(x - 4) - 8 you see how the function p(x) acts as x in the function k(x)?

OpenStudy (anonymous):

distribute first and then solve

OpenStudy (meowlover17):

ok

OpenStudy (meowlover17):

5x-28

OpenStudy (anonymous):

correct

OpenStudy (meowlover17):

Given the following functions f(x) and g(x), solve f over g (−5) and select the correct answer below: f(x) = 2x − 20 g(x) = x − 1

OpenStudy (meowlover17):

−5 5 one sixth 30

OpenStudy (meowlover17):

one sixth = 1/6

OpenStudy (meowlover17):

@090909090909

OpenStudy (anonymous):

5

OpenStudy (meowlover17):

Thank you :)

OpenStudy (anonymous):

i actually have to leave but if you have anymore questions open up a new thread and i'm sure you'll get answers

OpenStudy (meowlover17):

alright thanks :) have a great rest of your day